# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ B.2 The rate
In the rest frame of the incoming parton, $s=m_{Q}^{2}+2e^{\prime}_{q}m_{Q}+m_{q}^{2}\approx m_{Q}^{2}+2e^{\prime}_{q}m_{Q}$ and consequently $s_{=}\coloneqq s-s_{\rm min}=2e^{\prime}_{q}m_{Q}-m_{g}^{2}-2m_{g}m_{Q}$ and $\Theta(s-s_{\rm min})\Leftrightarrow e_{q}^{\prime}\geq\underbrace{m_{g}+\frac{m_{g}^{2}}{2m_{Q}}}_{e^{\prime}_{\rm min}}.$ Introducing the difference $\tilde{e}_{q}=e_{q}^{\prime}-\left(m_{g}+\frac{m_{g}^{2}}{2m_{Q}}\right)$ which ranges from 0 to +$\infty$+ ∞ $\Rightarrow s_{=}=2m_{Q}\tilde{e}_{q}$ and
@@
$=$ $\frac{2\pi\,n!}{u}\times\left(\frac{e^{-\frac{u_{Q}^{0}-u_{Q}}{T}e_{\rm min}^{\prime}}}{(u_{Q}^{0}-u)^{n+1}}-\frac{e^{-\frac{u_{Q}^{0}+u_{Q}}{T}e_{\rm min}^{\prime}}}{(u_{Q}^{0}+u_{Q})^{n+1}}\right)\times\left(\frac{T}{m_{Q}}\right)^{n+2}\,,$ (117)
where $u^{0}_{Q}=e_{Q}/m_{Q}$ and $u_{Q}=\|\vec{p}_{Q}\|/m_{Q}$, and finds
@@ B.3 The sampling process
    At this stage, we are dealing with accepted (pre)gluons and one just needs to generate the final kinematics. One recalls that $\vec{l}_{\perp}$ and $\vec{k}_{\perp}$ are interpreted as transverse momenta (only) in the frame where $\vec{p}_{Q}$ and $\vec{p}_{q}$ are collinear. In this frame, one should have the same energy before and after the collision $e^{\prime}_{q}=e_{q}$, for the parton of the QGP medium and hence also $e_{Q}=e^{\prime}_{Q}+\omega$. One can either generate the final kinematics in this frame and then boost back to the local QGP rest frame or evaluate invariant quantities that will be used to directly generate the final four-momenta in the local QGP frame. In the eikonal limit where $P^{+}$ is the largest scale and $k^{+}$ the second largest scale, the transferred four-momentum $l$ is dominated by its transverse component. Assuming $l^{+}=l^{-}=0$ however implies small violations of the energy-momentum conservation. These can be cured by an explicit evaluation of $l^{+}$ and $l^{-}$.
@@
    If one deals properly with the kinematics, one can demonstrate that $l^{+}$ and $l^{-}$ become of the same order as $l_{\perp}$ whenever $x\lesssim\frac{m_{g}^{2}+k_{\perp}^{2}}{{p_{Q}^{+}}^{2}}$, what corresponds to gluons radiated backwards. For this kinematic range, approximating the exchanged gluon propagator by $(l_{\perp}^{2}+\mu^{2})^{-1}$ as done in eq. (68) leads to a large excess of gluon radiated backwards. A pragmatic way to cure this problem — without building a new sampling strategy — consists in oversampling the initial rate by a factor ×$\times 2$ and permuting the role of $\vec{p}_{Q}$ and $\vec{p}_{q}$ once placed in the CM frame with $50\%$ probability (thus considering that the gluon is radiated ‘‘forward’’ by the the target parton in this frame) while rejecting gluons that are radiated in the opposite direction of their ‘‘radiator’’ ($y<0$ for gluons radiated by the projectile parton and $y>0$ for gluons radiated by the target parton).
Appendix C Sampling the Bethe-Heitler-Gunion-Bertsch seed for $m_{q}=+\infty$
We adapt the strategy provided in appendix B to the case of infinite mass, $m_{q}\to\infty$. Since this case is solely developed in order to perform the comparison with the BDMPS-Z benchmark results, we will assume an ultra-relativistic projectile and $x$ not too close to $1$.
C.1 The cross section
One starts from the definition (70), with $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})\to\Delta(x,\vec{0}_{\perp},\vec{0}_{\perp},m_{q})$ as an approximation of the measure. To derive the radiation cross section in the $m_{q}\to\infty$ limit, we naturally retain the leading terms of each factor, placing ourselves in the rest frame of the static scattering centers. $B^{+}$ (eq. (60)) can be reformulated as
@@ C.1 The cross section
One can convince oneself that the $B^{+}\geq 0$ condition simply expresses that $p^{\prime}_{Q,z}$, the projectile momentum along $Oz$ after the collision is real, using light-cone variables as defined in appendix A. [^32] Neglecting $p_{Q}^{-}$, $B^{+}>0\Leftrightarrow(1-x)\frac{p_{Q}^{+}}{2}-\frac{k^{-}}{2}>m_{Q,T}^{\prime}\Leftrightarrow e_{Q}--x\frac{p_{Q}^{+}}{2}-\frac{k^{-}}{2}>m_{Q,T}^{\prime}\Leftrightarrow e_{Q}-\omega>m_{Q,T}^{\prime}$, which is precisely the onshellness condition for the outgoing projectile. This translates, for a given $x$, into the condition for ${\vec{l}_{\perp},\vec{k}_{\perp}}$ in the transverse space
@@
$l_{\perp}^{\rm max}(k_{\perp})=\sqrt{\left(\frac{(1-x)p_{Q}^{+}-p_{Q}^{-}}{2}-\frac{m_{g}^{2}+k_{\perp}^{2}}{2xp_{Q}^{+}}\right)^{2}-m_{Q}^{2}}\,.$ (121)
@@
$k_{\perp}^{\rm max}\propto\sqrt{x}(1-\sqrt{x})p_{Q}^{+}\,,$ (122)
for an ultra-relativistic projectile.
As already announced, $\Delta_{\rm app}=\Delta(x,0,0,m_{q})$ – where the finite-$m_{q}$ expression of $\Delta$ is given in eq. (59) – is used in the expression of $\sigma_{\rm rad}$, with the $m_{q}\to\infty$ limit of $\Delta_{\rm app}$
$\Delta_{\rm app}\underset{m_{q}\to\infty}{=}(p_{Q}^{+})^{4}\left(x\left((1-x)+\frac{2m_{Q}}{p_{Q}^{+}}\right)-\frac{m_{g}^{2}}{{p_{Q}^{+}}^{2}}\right)\times\left(x\left((1-x)-\frac{2m_{Q}}{p_{Q}^{+}}\right)-\frac{m_{g}^{2}}{{p_{Q}^{+}}^{2}}\right)\,.$ (123)
At leading order in $\mathcal{O}(m_{g}/p_{Q}^{+})$ and $\mathcal{O}(m_{Q}/p_{Q}^{+})$, the $x$ boundaries $x^{\prime}_{\rm min}$ and $x^{\prime}_{\rm max}$ obtained by setting $\Delta_{\rm app}=0$ are $x^{\prime}_{\rm min}=\left(\frac{m_{g}}{p_{Q}^{+}}\right)^{2}$ and $x^{\prime}_{\rm max}=1-\frac{2m_{Q}}{p_{Q}^{+}}$. It should be noted that these definitions differ from the ones obtained in appendix B for finite $m_{q}$. This is the consequence of the non-commutativity between the $e_{Q}\to\infty$ and $m_{q}\to\infty$ limits. With these boundaries, one can write
$\sqrt{\Delta_{\rm app}}\approx(p_{Q}^{+})^{2}(x-x^{\prime}_{\rm min})\sqrt{(x^{\prime}_{\rm max}-x)(2-x^{\prime}_{\rm max}-x)}\approx(p_{Q}^{+})^{2}(x-x^{\prime}_{\rm min})(1-x)\,,$ (124)
where the $p_{Q}^{+}\to\infty$ limit was taken for $x^{\prime}_{\rm max}$, as will be assumed in the rest of this appendix. As $p_{\rm tot}^{+}\underset{m_{q}\to\infty}{\approx}m_{q}$ while $p_{Q}^{+}\approx 2e_{Q}$, one has
@@
Coming to the elastic amplitude (eq. (68)) one has $s_{-}=2m_{q}e_{Q}$ for a static quark $q$, as well as $v_{qQ}=m_{q}e_{Q}$ and thus
@@
Combining both eqs. (125) and (126), we obtain from eq. (70)
$\frac{\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}\to\infty}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}$$\approx$ $\frac{1}{32\pi^{2}}\frac{8C_{F}g^{4}}{N_{c}}\times\frac{1-\frac{l_{\perp}^{2}}{4e_{Q}^{2}}}{(l_{\perp}^{2}+\mu^{2})^{2}}\frac{x}{2(x-x^{\prime}_{\rm min})}P_{g}\times\Theta(B^{+})$ (127)
@@
$\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{1}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}<\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}$ $=$ $\frac{2C_{F}\alpha_{s}^{2}}{N_{c}(l_{\perp}^{2}+\mu^{2})^{2}}\times\frac{2C_{A}\alpha_{s}}{\pi}\frac{1-x}{x}\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)$ (130)
@@
as $C_{A}=N_{c}$. One should notice that this expression is quite similar to the one obtained in the $m_{q}\rightarrow 0$ limit — see eq. (104) — with $x^{\prime}_{\rm max}\to 1$ in the expression of $c_{2}(x)$. One can thus proceed as in the $m_{q}=0$ case, which ends with the definition of
@@
where $\tilde{m}_{g}^{\rm min}$ is here the smallest value of $\tilde{m}_{g}$ on $[x^{\prime}_{\rm min},x^{\prime}_{\rm max}]$. We find
$I_{2}\approx\ln\frac{1}{x^{\prime}_{\rm min}}=2\ln\frac{p_{Q}^{+}}{m_{g}}\,,$ (133)
@@
$\sigma_{\rm rad}^{{\rm sup}_{4}}\approx 8C_{F}\alpha_{s}^{3}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}-\mu^{2}}\ln\frac{p_{Q}^{+}}{m_{g}}\,.$ (134)
@@
C.2 The total radiative rate
@@
where $n_{q}$ should be considered as the density of partons (calculated in the $m_{q}=0$ approximation with a MB distribution [^33] We are indeed just probing the consequence of the $m_{q}=0$ vs $m_{q}=+\infty$ choice for the radiation, assuming the same density of scatters.):
@@ C.3 The sampling algorithm
    Generate the outgoing four-momentum of the outgoing gluon as $k^{+}=\omega+k_{L}=xp_{Q}^{+}=x(p_{Q}+e_{Q})\approx 2e_{Q}$ and $k^{-}=\frac{m_{g}^{2}+k_{\perp}^{2}}{k^{+}}\Rightarrow\omega=\frac{k^{+}+k^{-}}{2}=xe_{Q}+\frac{m_{g,T}^{2}}{4xe_{Q}}$.
@@
    Generate the outgoing four-momentum of the incoming parton following $p^{\prime}_{Q}=({p^{\prime}_{Q}}^{+}\coloneqq\tilde{p}^{+}_{Q}-k^{+},-\vec{k}_{\perp}+\vec{l}_{\perp},{p^{\prime}_{Q}}^{-}=\frac{m_{Q}^{2}+(-\vec{k}_{\perp}+\vec{l}_{\perp})^{2}}{{p^{\prime}_{Q}}^{+}})$.
Appendix D Calibration

### v2

@@ B.3.3 The rate for $m_{q}=0$
In the rest frame of the incoming parton $u_{f}=(u_{Q}^{0},\vec{u}_{Q})$, where $u_{Q}^{0}=e_{Q}/mQ$ and $\vec{u}_{Q}=\vec{p}_{Q}/m_{Q}$. Moreover $s=m_{Q}^{2}+2e^{\prime}_{q}m_{Q}+m_{q}^{2}\approx m_{Q}^{2}+2e^{\prime}_{q}m_{Q}$ and consequently $s_{=}\coloneqq s-s_{\rm min}=2e^{\prime}_{q}m_{Q}-m_{g}^{2}-2m_{g}m_{Q}$ and $\Theta(s-s_{\rm min})\Leftrightarrow e_{q}^{\prime}\geq\underbrace{m_{g}+\frac{m_{g}^{2}}{2m_{Q}}}_{e^{\prime}_{\rm min}}.$ Introducing the difference $\tilde{e}_{q}=e_{q}^{\prime}-\left(m_{g}+\frac{m_{g}^{2}}{2m_{Q}}\right)$ which ranges from 0 to +$\infty$+ ∞ $\Rightarrow s_{=}=2m_{Q}\tilde{e}_{q}$ and
@@
$=$ $\frac{2\pi\,n!}{u_{Q}}\times\left(\frac{e^{-\frac{u_{Q}^{0}-u_{Q}}{T}e_{\rm min}^{\prime}}}{(u_{Q}^{0}-u_{Q})^{n+1}}-\frac{e^{-\frac{u_{Q}^{0}+u_{Q}}{T}e_{\rm min}^{\prime}}}{(u_{Q}^{0}+u_{Q})^{n+1}}\right)\times\left(\frac{T}{m_{Q}}\right)^{n+2}\,,$ (123)
where $u_{Q}=\|\vec{u}_{Q}\|$, and finds
@@ B.3.4 The sampling process for $m_{q}=0$
    At this stage, we are dealing with accepted (pre)gluons and one just needs to generate the final kinematics assuming that the particles are on-shell. One recalls that $\vec{l}_{\perp}$ and $\vec{k}_{\perp}$ are interpreted as transverse momenta (only) in the frame where $\vec{p}_{Q}$ and $\vec{p}_{q}$ are collinear. One can either generate the final kinematics in this frame and then boost back to the local QGP rest frame or evaluate invariant quantities that will be used to directly generate the final four-momenta in the local QGP frame. In the eikonal limit where $p_{Q}^{+}$ is the largest scale and $k^{+}$ the second largest scale, the transferred four-momentum $l$ is dominated by its transverse component while $l^{+}$ and $l^{-}$ are both $\propto 1/p^{+}_{Q}$. Assuming $l^{+}=l^{-}=0$ however implies small violations of the energy-momentum conservation. These can be cured by an explicit evaluation of $l^{+}$ and $l^{-}$.
@@
    If one deals properly with the kinematics, one can demonstrate that $l^{+}$ and $l^{-}$ become of the same order as $l_{\perp}$ whenever $x\lesssim\frac{m_{g}^{2}+k_{\perp}^{2}}{{p_{Q}^{+}}^{2}}$, what corresponds to gluons radiated backwards. For this kinematic range, approximating the exchanged gluon propagator by $(l_{\perp}^{2}+\mu^{2})^{-1}$ as done in eq. (74) leads to a large excess of gluon radiated backwards. A pragmatic way to cure this problem — without building a new sampling strategy — consists in oversampling the initial rate by a factor ×$\times 2$ and permuting the role of $\vec{p}_{Q}$ and $\vec{p}_{q}$ once placed in the CM frame with $50\%$ probability (thus considering that the gluon is radiated ‘‘forward’’ by the target parton in this frame) while rejecting gluons that are radiated in the opposite direction of their ‘‘radiator’’ ($y<0$ for gluons radiated by the projectile parton and $y>0$ for gluons radiated by the target parton).
B.4 Sampling the Bethe-Heitler-Gunion-Bertsch seed for $m_{q}=+\infty$
We now adapt the strategy to the case of infinite mass, $m_{q}\to\infty$. Since this case is solely developed in order to perform the comparison with the BDMPS-Z benchmark results, we will assume an ultra-relativistic projectile and $x$ not too close to $1$.
One starts from the definition (76), with $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})\to\Delta(\tilde{x},\vec{0}_{\perp},\vec{0}_{\perp},m_{q})$ as a good approximation of the measure, as argued previously in this Appendix: B.3.1.
B.4.1 Phase space for $m_{q}\to\infty$
To derive the phase space and the radiation cross section in the $m_{q}\to\infty$ limit, we naturally retain the leading terms of each factor, placing ourselves in the rest frame of the static scattering centers. $B^{+}$ (eq. (61)) can be reformulated as
@@ B.4.1 Phase space for $m_{q}\to\infty$
One can convince oneself that the $B^{+}\geq 0$ condition simply expresses that $p^{\prime}_{Q,z}$, the projectile momentum along $Oz$ after the collision is real, using light-cone variables as defined in appendix A. [^32] Neglecting $p_{Q}^{-}$, $B^{+}>0\Leftrightarrow(1-x)\frac{p_{Q}^{+}}{2}-\frac{k^{-}}{2}>m_{Q,T}^{\prime}\Leftrightarrow e_{Q}-x\frac{p_{Q}^{+}}{2}-\frac{k^{-}}{2}>m_{Q,T}^{\prime}\Leftrightarrow e_{Q}-\omega>m_{Q,T}^{\prime}$, which is precisely the onshellness condition for the outgoing projectile. This translates, for a given $x$, into the condition for ${\vec{l}_{\perp},\vec{k}_{\perp}}$ in the transverse space
@@
$l_{\perp}^{\rm max}(k_{\perp})=\sqrt{\left(\frac{(1-x)p_{Q}^{+}+p_{Q}^{-}}{2}-\frac{m_{g}^{2}+k_{\perp}^{2}}{2xp_{Q}^{+}}\right)^{2}-m_{Q}^{2}}\,.$ (127)
@@
$k_{\perp}^{\rm max}\propto\sqrt{x}\sqrt{1-x}p_{Q}^{+}\,,$ (128)
for an ultra-relativistic projectile. As already announced, $\Delta_{\rm a2}=\Delta(x,0,0,m_{q})$ – where the finite-$m_{q}$ expression of $\Delta$ is given in eq. (60) – is used in the expression of $\sigma_{\rm rad}$, with the $m_{q}\to\infty$ limit of $\Delta_{\rm a2}$
$\Delta_{\rm a2}\underset{m_{q}\to\infty}{=}(p_{Q}^{+})^{4}\left(x\left((1-x)+\frac{2m_{Q}}{p_{Q}^{+}}\right)-\frac{m_{g}^{2}}{{p_{Q}^{+}}^{2}}\right)\times\left(x\left((1-x)-\frac{2m_{Q}}{p_{Q}^{+}}\right)-\frac{m_{g}^{2}}{{p_{Q}^{+}}^{2}}\right)\,.$ (129)
At leading order in $\mathcal{O}(m_{g}/p_{Q}^{+})$ and $\mathcal{O}(m_{Q}/p_{Q}^{+})$, the $x$ boundaries $x^{\prime}_{\rm min}$ and $x^{\prime}_{\rm max}$ obtained by setting $\Delta_{\rm a2}=0$ are
$x^{\prime}_{\rm min}=\left(\frac{m_{g}}{p_{Q}^{+}}\right)^{2}\quad\text{and}\quad x^{\prime}_{\rm max}=1-\frac{2m_{Q}}{p_{Q}^{+}}.$ (130)
It should be noted that these definitions differ from the ones obtained previously for finite $m_{q}$. This is the consequence of the non-commutativity between the $e_{Q}\to\infty$ and $m_{q}\to\infty$ limits. With these boundaries, one can write
$\sqrt{\Delta_{\rm a2}}\approx(p_{Q}^{+})^{2}(x-x^{\prime}_{\rm min})\sqrt{(x^{\prime}_{\rm max}-x)(2-x^{\prime}_{\rm max}-x)}\approx(p_{Q}^{+})^{2}(x-x^{\prime}_{\rm min})(1-x)\,,$ (131)
where the $p_{Q}^{+}\to\infty$ limit was taken for $x^{\prime}_{\rm max}$, as will be assumed in the rest of this appendix.
B.4.2 Explicit radiative cross section for $m_{q}\to\infty$
As $p_{\rm tot}^{+}\underset{m_{q}\to\infty}{\approx}m_{q}$ while $p_{Q}^{+}\approx 2e_{Q}$, one has
@@ B.4.2 Explicit radiative cross section for $m_{q}\to\infty$
Injecting this result in eq. (76) yields
$\frac{\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}\to\infty}}{\mathrm{d}x\mathrm{d}^{2}l_{\perp}\mathrm{d}^{2}}k_{\perp}\approx\frac{\pi^{3}\left|\mathcal{M}_{\rm el}\right|^{2}}{(2\pi)^{5}v_{qQ}}\frac{1}{2m_{q}e_{Q}}\frac{x}{(x-x^{\prime}_{\rm min})}P_{g}\Theta(B^{+})\,,$ (133)
an expression which is formally similar to eq. 101 for the $m_{q}=0$ case. Coming to the elastic amplitude (eq. (74)) one has $s_{-}=2m_{q}e_{Q}$ for a static quark $q$, as well as $v_{qQ}=m_{q}e_{Q}$ and thus
@@
Combining both eqs. (133) and (134), we obtain
$\frac{\mathrm{d}^{5}\sigma_{\rm rad}^{m_{q}\to\infty}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}$$\approx$ $\frac{1}{64\pi^{2}}\frac{8C_{F}g^{4}}{N_{c}}\times\frac{1-\frac{l_{\perp}^{2}}{4e_{Q}^{2}}}{(l_{\perp}^{2}+\mu^{2})^{2}}\frac{x}{(x-x^{\prime}_{\rm min})}P_{g}\times\Theta(B^{+})$ (135)
@@
$\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{1}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}<\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}$ $=$ $\frac{2C_{F}\alpha_{s}^{2}}{N_{c}(l_{\perp}^{2}+\mu^{2})^{2}}\times\frac{2C_{A}\alpha_{s}}{\pi}\frac{1-x}{x-x^{\prime}_{\rm min}}\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)$ (138)
@@
as $C_{A}=N_{c}$. One should notice that this expression is quite similar to the one obtained in the $m_{q}\rightarrow 0$ limit — see eq. (110) — with $x^{\prime}_{\rm max}\to 1$ and $x_{\rm min}\to x^{\prime}_{\rm min}$ in the expression of $c_{2}(x)$. One can thus proceed as in the $m_{q}=0$ case, which ends with the definition of
@@
where $\tilde{m}_{g}^{\rm min}$ is here the smallest value of $\tilde{m}_{g}$ on $[x^{\prime}_{\rm min},1]$. We find
$I_{2}\approx\ln\frac{1}{x^{\prime}_{\rm min}}=\ln\frac{(p_{Q}^{+})^{2}}{m_{g}^{2}}\,,$ (141)
@@
$\sigma_{\rm rad}^{{\rm sup}_{4}}\approx 4C_{F}\alpha_{s}^{3}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}-\mu^{2}}\ln\frac{(p_{Q}^{+})^{2}}{m_{g}^{2}}\,.$ (142)
Comparing this expression with the equivalent one for the $m_{q}=0$ case (eq. 114) we see that for the $m\to\infty$ limit s is replaced by $(P_{Q}^{+})^{2}$, a relation which we found already when comparing the $k_{\perp}^{\rm max}$ values for the $m_{q}=0$ (eq. 85) and the $m_{q}\to\infty$ (eq. 128) case.
@@
B.4.3 The total radiative rate for $m_{q}\to\infty$
@@
where $n_{q}$ should be considered as the density of partons (calculated in the $m_{q}=0$ approximation with a MB distribution [^33] We are indeed just probing the consequence of the $m_{q}=0$ vs $m_{q}=+\infty$ choice for the radiation, assuming the same density of scattering centers.):
@@ B.4.4 The sampling algorithm for $m_{q}\to\infty$
    Check whether the sampled variables satisfy the boundary condition $B^{+}>0$. If not, reject the radiation candidate.
@@
    Generate the four-momentum of the outgoing gluon as $k^{+}=\omega+k_{z}=xp_{Q}^{+}=x(p_{Q}+e_{Q})\approx 2e_{Q}$ and $k^{-}=\frac{m_{g}^{2}+k_{\perp}^{2}}{k^{+}}\Rightarrow\omega=\frac{k^{+}+k^{-}}{2}=xe_{Q}+\frac{m_{g,T}^{2}}{4xe_{Q}}$.
@@
    Generate the outgoing four-momentum of the parton applying energy conservation ($e^{\prime}_{Q}=e_{Q}-\omega$), while $\vec{p}\,^{\prime}_{Q\perp}=\vec{l}_{\perp}-\vec{k}_{\perp}$ and $p^{\prime}_{Qz}$ follows from requiring that the parton is on-shell. Let us recall that in our MC we have two options to shift from $p_{Q}\to p^{\prime}_{Q}=p_{Q}+l-k$: This shift can be performed at the time when the virtual gluon is emitted or when it is effectively formed as a real gluon, see discussion in section IV.5.2.
Appendix C Calibration